For the points P and Q, find the distance d(P,Q)
P(-5, -7), Q(-13,1)

Found 2 solutions by user_dude2008, nyc_function:
Answer by user_dude2008(1862)

d(P,Q) = sqrt((-5+13)^2+(-7-1)^2) = sqrt(8^2+(-8)^2) = sqrt(64+64) = sqrt(128) = 8*sqrt(2) = 11.3137

Answer by nyc_function(2741)

Let d = distance from P to Q.
d = sqrt{(-13 + 5)^2 + (1 + 7)^2}
d = sqrt{(-8)^2 + (8)^2}
d = sqrt{64 + 64}
d = sqrt{128}
d = sqrt{16} * sqrt{8}
d = 4 * sqrt{4} * sqrt{2}
d = 4 * 2 * sqrt{2}
d = 8(sqrt{2})